Macho14:

Math help please. I know this is a system of equations, but am not sure to go about how setting the problem up.

Macho14:

1 attachment
Macho14:

This method is incorrect. I wrote in the question that this is a system of equations.

surjithayer:

let x ml of 70 % alcohol be added to 45 ml of 10 % alcohol to make (x+45) ml of 65 % alcohol. 70 %=0.7 10%=0.1 65%=0.65 0.7x+0.1*45=0.65(x+45) multiply by 100 70x+450=65(x+45) 70x+450=65x+65 *45 70x-65x=65*45-450 5x=45(65-10) 5x=65*55 \[x=\frac{ 45\times 55 }{ 5 }=45\times11=495~ml\]

Macho14:

thanks

surjithayer:

yw